Which of the following instruments pays the holder of the instrument a fixed interest payment every year until maturity, and then pays the holder the face value (principle) of the instrument? at maturity O A. fixed-payment loan O B. simple loan O C. coupon bond O D. discount bond Suppose that a bond has one year to maturity. The yield to maturity on the bond if it was bought for $1130.00 and has a $1100 face value with a coupon rate of 11% Round your response to the nearest whole number Consider a coupon bond with a face value of S 1050, one year to maturity and a coupon rate of 6% Given a yield to maturity of 5%, the price the bond will sll for is s (Round your response to the nearest two decimal place)
Answer #1

(1) Option (C)

Coupon bond pays an annual coupon every year and pays back the face value at end of maturity.

(2) Yield to maturity (YTM) = (Bond face value + Annual coupon - Purchase price) / Purchase price

Annual coupon = $1100 x 11% = $121

YTM = $(1100 + 121 - 1130) / $1130 = $91 / $1130 = 0.08% = 8%

(3) Bond price = Present value of annual coupon + Present value of face value

Annual coupon = $1050 x 6% = $63

Bond price ($) = 63 x P/F(5%, 1) + 1050 x P/F(5%, 1) = (63 + 1050) x P/F(5%, 1) = 1113 x 0.9524** = 1060.02

**From P/F Factor table